# Exploiting Spatial Diversity in Earth-to-Satellite Quantum-Classical Communications

### v1

@@ 4 Log-Normal Model for Channel Loss
ln(Loss_LgN) ∼N (${{\hbox{\operatorname{ln}({\text{Loss}}_{\text{LgN}}) \sim\mathcal{N} \left(\ln\left(\frac{\mu_{\text{LgN}}^{2}}{\sqrt{\mu_{\text{LgN}}^{2}+\sigma_{\text{LgN}}^{2}}}\right),\ln\left(1+\frac{\sigma_{\text{LgN}}^{2}}{\mu_{\text{LgN}}^{2}}\right)\right)}}},$ (A5)
@@
Appendix B Diversity Combining
In an $M\times M$ diversity scheme using $M$ subchannels, at least $M-1$ beam splitters are required in a combining module to combine the $M$ incoming signals. We denote the $M$ incoming quantum states to be combined as $\smash{\{\rho_{j}\}}$ ($\smash{1\leq j\leq M}$), the transmission coefficient of the $l^{\text{th}}$ beam splitter ($\smash{1\leq l\leq M-1}$) as $\eta_{l}$ ($\smash{0\leq\eta_{l}\leq 1,\,\,\eta_{l}\in\mathbb{R}}$), and the resulting quantum state from the combining of the $M$ quantum states as $\rho_{\text{DC}}^{(M)}$.
@@ Appendix B Diversity Combining
Appendix C Analytical Expression of Fidelity
Denoting the encoded coherent state as $|\alpha_{\text{tar}}\rangle$ and the transmitted coherent state (same for all the transmitters due to the equal splitting) as $|\alpha_{\text{tx}}\rangle=|\frac{\alpha_{\text{tar}}}{\sqrt{M}}\rangle$, we can analytically express the fidelity (before being averaged over the quantum alphabet and the fluctuating subchannel losses) in Eq. (22) as
@@ Appendix C Analytical Expression of Fidelity
where $|\alpha_{\text{tar}}^{\prime}\rangle=|{\alpha_{\text{tar}}}/{\langle\sqrt{T}\rangle}\rangle$ denotes the scaled encoded state (recall Sec. IV.2). In Eq. (A6), $X_{(M)}$ and $Y_{(M)}$ are given by
$X_{(M)}=\begin{cases}\sqrt{T_{1}},\,\,&M=1\\ \eta_{1}\sqrt{T_{1}}+\sqrt{1-\eta_{1}^{2}}\sqrt{T_{2}},\,\,&M=2\\ \eta_{2}\left(\eta_{1}\sqrt{T_{1}}+\sqrt{(1-\eta_{1}^{2})T_{2}}\right)+\sqrt{(1-\eta_{2}^{2})T_{3}},\,\,&M=3\\ \eta_{3}\left(\eta_{1}\sqrt{T_{1}}+\sqrt{(1-\eta_{1}^{2})T_{2}}\right)+\eta_{2}\sqrt{(1-\eta_{3}^{2})T_{3}}+\sqrt{(1-\eta_{2}^{2})(1-\eta_{3}^{2})T_{4}},\,\,&M=4\end{cases}$ (A7)
@@
where $\{T_{j}\}$ denote the subchannel transmissivities ($1\leq j\leq M$), $\{\epsilon_{j}\}$ denote the subchannel excess noises ($1\leq j\leq M$), and $\{\eta_{l}\}$ denotes the transmission coefficients of the combining beam splitters ($11\leq l\leq M-$ 1). For generality, in Eqs. (A7)(A8) we have chosen not to plug in any specific values for $\{\eta_{l}\}$.
@@
Section S-III Beam Splitter Transformation
A General Description
As a Gaussian transformation that transforms Gaussian states into Gaussian states, the beam splitter transformation is one of the most important linear optics transformations for two bosonic modes. The convention of the beam splitter transformation adopted in this work is underpinned by the beam splitter model illustrated in Fig. S1. Specifically, a beam splitter with a real transmission coefficient $\eta$ ($0\leq\eta\leq 1$, $\eta\in\mathbb{R}$) performs the following linear mode transformation in the Heisenberg picture
$\begin{bmatrix}\hat{a}_{\text{out,1}}\\ \hat{a}_{\text{out,2}}\end{bmatrix}=\begin{bmatrix}\eta&\sqrt{1-{\eta}^{2}}\\ \sqrt{1-{\eta}^{2}}&-\eta\end{bmatrix}\begin{bmatrix}\hat{a}_{\text{in,1}}\\ \hat{a}_{\text{in,2}}\end{bmatrix},$ (S13)
where $\hat{a}_{\text{in,1}}$, $\hat{a}_{\text{in,2}}$, $\hat{a}_{\text{out,1}}$, and $\hat{a}_{\text{out,2}}$ denote the annihilation operators for input mode 1, input mode 2, output mode 1 (i.e., the transmitted mode), and output mode 2 (i.e., the reflected mode), respectively.
Figure S1: The beam splitter model adopted in this work.
Under the CF formalism, the beam splitter transformation in Eq. (S13) corresponds to a CF argument transformation that leads to the following result
$\chi_{\text{out}}(\xi_{1},\xi_{2})=\chi_{\text{in}}\left(\eta\xi_{1}+\sqrt{1-{\eta}^{2}}\xi_{2},\sqrt{1-{\eta}^{2}}\xi_{1}-\eta\xi_{2}\right),$ (S14)
$\chi_{\text{out},1}(\xi_{1})=\chi_{\text{out}}(\xi_{1},0)=\chi_{\text{in}}\left(\eta\xi_{1},\sqrt{1-{\eta}^{2}}\xi_{1}\right).$ (S15)
The statistical-moment description can also be conveniently used to describe a beam splitter transformation. Specifically, for the linear beam-splitter transformation given in Eq. (S13), the quadrature operators $\hat{\mathbf{x}}_{\text{in}}=\left[\hat{q}_{\text{in,1}},\hat{p}_{\text{in,1}},\hat{q}_{\text{in,2}},\hat{p}_{\text{in,2}}\right]^{\top}$ are transformed via the symplectic map
$\hat{\mathbf{x}}_{\text{out}}=\mathbf{B}(\eta)\hat{\mathbf{x}}_{\text{in}},$ (S16)
$\bar{\mathbf{x}}_{\text{out}}=\mathbf{B}(\eta)\bar{\mathbf{x}}_{\text{in}},\quad\mathbf{V}_{\text{out}}=\mathbf{B}(\eta)\mathbf{V}_{\text{in}}\mathbf{B}(\eta)^{\top}.$ (S17)
B Beam-Splitter Description of Atmospheric Channel
The beam splitter transformation is ubiquitously used in this work. Specifically, we follow the common practice and model an atmospheric quantum channel as a lossy channel with a real transmissivity $T$ and a total excess noise $\epsilon$ (see e.g., [1, 2]). Such a channel preserves the Gaussian characters of a quantum state and can be conveniently modelled as a beam splitter in a thermal background with a certain number of mean photons. In this work, we explicitly describe an atmospheric quantum channel as a beam splitter with input mode $\hat{a}_{\text{in,1}}$ being the quantum signal and input mode $\hat{a}_{\text{in,2}}$ being the associated environmental mode in a thermal state with equal quadrature variance $V_{\text{thermal}}=1+\frac{\epsilon}{1-T}$ [2].
C Beam-Splitter Description of Diversity Combining

### v2

@@ 2 Phase-Screen Simulations for Realistic Uplink Channels
Despite being computationally intensive, our phase-screen simulations represent one of the most powerful and flexible methods to investigate optical propagation within the atmosphere of the Earth. Although our simulations do not yield analytical results, they implicitly capture all the well-known turbulence-induced effects (including scintillation, beam wandering, beam deformation, and beam broadening) without relying on the approximations (e.g., those related to certain turbulence parameters and the final beam shape) commonly employed by state-of-the-art theoretical studies (e.g., [56, 15, 57, 58, 59]). Our simulations have been extensively applied within the context of terrestrial and satellite-based optical quantum communications (e.g., [5, 19, 54, 55, 60]), with their accuracy confirmed through real-world experiments [60] and cross-validated in [19] against state-of-the-art analytical models, particularly the family of models based on the elliptic-beam approximation (originally proposed in [56]). We refer the reader to [19] for a review of the literature covering other state-of-the-art channel modeling techniques.
@@ 4 Log-Normal Model for Channel Loss
$\operatorname{ln}({\text{Loss}}_{\text{LgN}})\sim\mathcal{N}\left(\ln\left(\frac{\mu_{\text{LgN}}^{2}}{\sqrt{\mu_{\text{LgN}}^{2}+\sigma_{\text{LgN}}^{2}}}\right),\ln\left(1+\frac{\sigma_{\text{LgN}}^{2}}{\mu_{\text{LgN}}^{2}}\right)\right),$ (A5)
@@
Appendix B Beam Splitter Transformation
1 General Description
As a Gaussian transformation that transforms Gaussian states into Gaussian states, the beam splitter transformation is one of the most important linear optics transformations for two bosonic modes. The convention of the beam splitter transformation adopted in this work is underpinned by the beam splitter model illustrated in Fig. A2. Specifically, a beam splitter with a real transmission coefficient $\eta$ ($0\leq\eta\leq 1$, $\eta\in\mathbb{R}$) performs the following linear mode transformation in the Heisenberg picture
$\begin{bmatrix}\hat{a}_{\text{out,1}}\\ \hat{a}_{\text{out,2}}\end{bmatrix}=\begin{bmatrix}\eta&\sqrt{1-{\eta}^{2}}\\ \sqrt{1-{\eta}^{2}}&-\eta\end{bmatrix}\begin{bmatrix}\hat{a}_{\text{in,1}}\\ \hat{a}_{\text{in,2}}\end{bmatrix},$ (A6)
where $\hat{a}_{\text{in,1}}$, $\hat{a}_{\text{in,2}}$, $\hat{a}_{\text{out,1}}$, and $\hat{a}_{\text{out,2}}$ denote the annihilation operators for input mode 1, input mode 2, output mode 1 (i.e., the transmitted mode), and output mode 2 (i.e., the reflected mode), respectively. Note that we reserve the use of the symbol $\eta$ (without any subscript) exclusively for the transmission coefficient in the general description of the beam splitter transformation.
Figure A2: The beam splitter model adopted in this work.
Under the CF formalism, the beam splitter transformation in Eq. (A6) corresponds to a CF argument transformation that leads to the following result
$\chi_{\text{out}}(\xi_{1},\xi_{2})=\chi_{\text{in}}\left(\eta\xi_{1}+\sqrt{1-{\eta}^{2}}\xi_{2},\sqrt{1-{\eta}^{2}}\xi_{1}-\eta\xi_{2}\right),$ (A7)
$\chi_{\text{out},1}(\xi_{1})=\chi_{\text{out}}(\xi_{1},0)=\chi_{\text{in}}\left(\eta\xi_{1},\sqrt{1-{\eta}^{2}}\xi_{1}\right).$ (A8)
The statistical-moment description can also be conveniently used to describe a beam splitter transformation. Specifically, for the linear beam-splitter transformation given in Eq. (A6), the quadrature operators $\hat{\mathbf{x}}_{\text{in}}=\left[\hat{q}_{\text{in,1}},\hat{p}_{\text{in,1}},\hat{q}_{\text{in,2}},\hat{p}_{\text{in,2}}\right]^{\top}$ are transformed via the symplectic map
$\hat{\mathbf{x}}_{\text{out}}=\mathbf{B}(\eta)\hat{\mathbf{x}}_{\text{in}},$ (A9)
$\bar{\mathbf{x}}_{\text{out}}=\mathbf{B}(\eta)\bar{\mathbf{x}}_{\text{in}},\quad\mathbf{V}_{\text{out}}=\mathbf{B}(\eta)\mathbf{V}_{\text{in}}\mathbf{B}(\eta)^{\top}.$ (A10)
2 Beam-Splitter Description of Atmospheric Channel
The beam splitter transformation is ubiquitously used in this work. Specifically, we follow the common practice and model an atmospheric quantum channel as a lossy channel with a real transmissivity $T$ and a total excess noise $\epsilon$ (see e.g., [49, 36]). Such a channel preserves the Gaussian characters of a quantum state and can be conveniently modelled as a beam splitter (with transmission coefficient $\sqrt{T}$) in a thermal background with a certain mean number of photons. In this work, we explicitly describe an atmospheric quantum channel as a beam splitter with input mode $\hat{a}_{\text{in,1}}$ being the quantum signal and input mode $\hat{a}_{\text{in,2}}$ being the associated environmental mode in a thermal state with equal quadrature variance $V_{\text{thermal}}=1+\frac{\epsilon}{1-T}$ [36].
3 Beam-Splitter Description of Diversity Combining
Appendix C Diversity Combining
In an $M\times M$ diversity scheme using $M$ subchannels, at least $M-1$ beam splitters are required in a combining module to combine the $M$ incoming signals. We denote the $M$ incoming quantum states to be combined as $\smash{\{\rho_{j}\}}$ ($\smash{1\leq j\leq M}$), the transmission coefficient of the $l^{\text{th}}$ combining beam splitter ($\smash{1\leq l\leq M-1}$) as $\eta_{l}$ ($\smash{0\leq\eta_{l}\leq 1,\,\,\eta_{l}\in\mathbb{R}}$), and the resulting quantum state from the combining of the $M$ quantum states as $\rho_{\text{DC}}^{(M)}$.
@@ Appendix C Diversity Combining
Here we elaborate on the real-world practicality of using the combining module, which consists mainly of beam splitters (note that auxiliary phase shifters may be needed in practice). The beam splitter is one of the simplest optical devices required in every quantum or classical optical communications system. In the standard practical implementation of free-space optical communications where the signal beam illuminating a tracked receiving aperture is first coupled into a fiber, the combining module can be constructed using multiple fiber-based beam splitters, which are virtually identical to their free-space counterparts. The combing module can also be integrated entirely onto a photonic chip (see an example of a 15-subchannel coherent combiner for FSO communications in [63]), further simplifying the combining operation.
Appendix D Analytical Expression of Fidelity
Denoting the encoded coherent state as $|\alpha_{\text{tar}}\rangle$ and the transmitted coherent state (same for all the transmitters due to the equal splitting) as $|\alpha^{(M)}_{\text{tx}}\rangle=|\frac{\alpha_{\text{tar}}}{\sqrt{M}}\rangle$, we can analytically express the fidelity (before being averaged over the quantum alphabet and the fluctuating subchannel losses) in Eq. (22) of the main text as
@@ Appendix D Analytical Expression of Fidelity
where $|\alpha_{\text{tar}}^{\prime}\rangle=|{\alpha_{\text{tar}}}/{\langle\sqrt{T}\rangle}\rangle$ denotes the scaled encoded state (recall Sec. IV.2 of the main text). In Eq. (A11), $X_{(M)}$ and $Y_{(M)}$ are given by
$X_{(M)}=\begin{cases}\sqrt{T_{1}},\,\,&M=1\\ \eta_{1}\sqrt{T_{1}}+\sqrt{1-\eta_{1}^{2}}\sqrt{T_{2}},\,\,&M=2\\ \eta_{2}\left(\eta_{1}\sqrt{T_{1}}+\sqrt{(1-\eta_{1}^{2})T_{2}}\right)+\sqrt{(1-\eta_{2}^{2})T_{3}},\,\,&M=3\\ \eta_{3}\left(\eta_{1}\sqrt{T_{1}}+\sqrt{(1-\eta_{1}^{2})T_{2}}\right)+\eta_{2}\sqrt{(1-\eta_{3}^{2})T_{3}}+\sqrt{(1-\eta_{2}^{2})(1-\eta_{3}^{2})T_{4}},\,\,&M=4\end{cases},$ (A12)
@@
where $\{T_{j}\}$ denote the subchannel transmissivities ($1\leq j\leq M$), $\{\epsilon_{j}\}$ denote the subchannel excess noises ($1\leq j\leq M$), and $\{\eta_{l}\}$ denote the transmission coefficients of the combining beam splitters ($11\leq l\leq M-$ 1). For generality, in Eqs. (A12)(A13) we have chosen not to plug in any specific values for $\{\eta_{l}\}$.
@@
Section S-III Usefulness of Uplink Configuration in Quantum Communications
In this section, we extend our discussion on the use of the uplink configuration for quantum communications and elaborate on its potential (and sometimes overlooked) usefulness within the context of current technological reach.
One may argue that the use of uplink channels should be entirely avoided since it is more challenging to transfer quantum signals in the uplink than in the downlink channels, and some of the well-known quantum information protocols can be implemented using solely downlink channels. Indeed, the original GG02 CV-QKD protocol [3] can arguably be implemented without transferring any quantum signal in the uplink, and in its entanglement-based implementation, the QKD protocol can be carried out using the shared entanglement established in either uplink or downlink. However, it remains unclear whether the need for uplink channels can be completely eliminated when other quantum information protocols are used in a practical deployment. For example, many advanced QKD protocols, such as measurement-device-independent QKD (e.g., [4, 5]), twin-field QKD (e.g.,  [6, 7]), and more recently proposed protocols (e.g., [8]), may naturally require the use of two uplink channels and a satellite to support the secret key generation between two distanced ground stations. For security reasons, using the uplink configuration also makes the attacks directed at receivers significantly more difficult in practice (see discussions in, e.g., [9]).
In addition, using the uplink configuration avoids deploying high-quality CV quantum entanglement sources on satellites, which is still impractical from a system complexity point of view. Currently, most space-ready quantum entanglement sources (e.g., [10]) generate polarization-entangled photon pairs, which are incompatible with CV quantum information protocols. A high-quality CV quantum entanglement source can take up considerable space in the payload and have stringent requirements such as ultra-stable laser light and fine-tuning of phase-matching achieved through precise adjustment and feedback control of the crystal temperature. The technical details revealing the high complexity of a high-quality squeezed laser source can be found in, e.g., [11] – it should be noted that one standard method for generating strong and stable CV entanglement requires two squeezed laser sources like this (see, e.g., [12]). The on-chip generation of squeezing via nanophotonics represents one promising direction towards a space-ready CV quantum entanglement source; however, despite the recent advancements (e.g., [13, 14]), the on-chip generation of strong squeezing has yet to be demonstrated, let alone a space-ready CV entanglement source. The use of the uplink configuration eliminates the need for (potentially complex) quantum entanglement sources in space, minimizing the complexity of the satellite and its payload. Such a minimization will be translated to reduced risk and cost as far as a practical deployment is concerned.
From an operational point of view, using the uplink configuration circumvents the significant limitations in flexibility and versatility of the downlink configuration. Despite advancements in aerospace technologies, upgrading, reconfiguring, diagnosing, or even regularly maintaining a quantum light source (especially a CV entanglement source) remains a challenging task in the downlink configuration (where the source is onboard an in-orbit satellite). On the contrary, using the uplink configuration benefits from deploying quantum light sources at the ground stations, bringing great flexibility and allowing for future upgrades. The satellite payload can also be more versatile (than one containing the quantum light source) since it may naturally be compatible with various quantum information protocols, including those that simply require the satellite to play the role of a relay. In principle, a satellite-based system adopting the uplink configuration can support various quantum information protocols using different quantum light sources since the quantum light source can be readily reconfigured or switched out at the ground stations.
Despite the higher loss in the uplink configuration, it can be argued that deploying a high-quality CV quantum source is still much more challenging than fitting a sufficiently large telescope onto a satellite. In fact, the Micius experiment [15] has demonstrated that the satellite-borne telescope aperture for uplink quantum communications does not have to be very large. We also note that even for downlink quantum communications, the satellite still needs a sufficiently large telescope to ensure a good beam quality (and to reduce diffraction). Furthermore, it can be expected that the future adoption of advanced quantum information protocols (e.g., [6, 7, 8]) will allow for a higher loss tolerance, fully exploiting the advantages provided by the use of the uplink configuration.